From GlueXWiki
Jump to: navigation, search

April 26, 2012 FDC meeting


  1. Final results from the oxygen studies (Lubomir)
  2. Production Construction Tracking (Dave)
    • Production status
    • Third package refurbishment status
    • Third package tests (Lubomir)
  3. Engineering update (Bill)
  4. Electronics update (Chris)
  5. Tests with first package at 126 (Beni)
    • Grounding
  6. Other


Participants: Fernando, Bill, Dave, Chris, Nick, Simon, Beni, Eugene, Glenn, and Lubomir.

Final results from the oxygen studies

- One full production cell with Viton O-rings and Apiezon applied on one side showed 75ppm (after 5 days flushing with gas) at 220ccpm gas flow. Bill got information from Servomex, the oxygen sensor manufacturer, that the flow defines the reaction time but effect on the reading is very small. With the same configuration at 110ccpm we had 114ppm, and with a flow of 55ccpm - 195ppm. These numbers can be explained if assuming the gas has 35ppm and the chamber contributes 40ppm at 220ccpm, therefore 80ppm at 110ccpm (80+35=115ppm) and 160ppm at 55ccpm (160+35=195ppm).


- Dave: for package #3 still two wire planes remain to be deadened, will do it tomorrow. Package #4: working on different cathodes, stringing third wire frame. Starting from the next wire frame we will first coat the G10 groove with Hysol and then do the stringing. The last wire frame #18, that required additional machining of the holes will be back from machine shop tomorrow.

- Two more cells were added to package #3 without tightening the package and tomorrow we will add another one, will have four in total. Then we will start flushing with gas and testing all four together. Normally we do the installation/testing cell by cell, but we want to save time on the flushing (5 days) and minimize the disassembling of the top end window and type-3 cathode that we have to do before each new cell installation. Bill: we take some risk with this, at least we can test the gas tightness by putting Lexan sheet on the top (after each cell installation). We will decide about this based on the experience with the first four cells. Eugene: when we can expect the third package ready for tests; if everything is OK we will need two weeks flushing and testing the first four cells, then another 1.5 weeks for the remaining two cells and one more week for HV cabling, pre-amps and cooling tubes and grounding.

- The first cell in package #3 was tested with a source. With low oxygen, the chamber is much more quite, very low noise (3-5mV), signals are higher and very uniform, and the working HV can be as low as 2050V where the chamber was stable for ~24 hours. Above that HV it was not possible to have it stable for a long time, the adverse effect of the low oxygen, which was helping before as a quencher. Also with the Viton the gas tightness is not as good as with EPDM: we have a leak (10^-4-10^-3 mL/sec) most likely due to a deformation in the gusset ring at that place. We found with Viton it's better to apply 100 in*oz torque on the rods, that's why we need them aluminum. Fernando pointed out that in this case we need to insulate them since we are closing the loops on the gusset rings. For that, we decided we better put the cuts on the top and bottom gusset rings at the same place.


- Bill is working on the fiducial scheme, spacers between the packages and on the tooling to remove from a package a whole cell together.

- After the meeting Bill will try to shim the place at the gusset ring causing the gas leakage. Also when installing the last cell of the third package, one of the metal pins didn't go through. Bill will figure out if the pin is bent or (if needed) we can ream the holes in the frames as installed in the package.


- We will get 5 more pre-amp sets to be used to ground the strips when testing package #3.

- Fernando: tests with the new fADC125 are going very well. Expect to be ready by June, at the same time working on the PR for the mass production. We expect to get two (out three) of them, so together with the old fADC125 we will have 3 for testing the packages at 126. We received first article LV cables, after minor corrections, the mass production will start soon. By June we can have the LV system ready to be used for the package tests in 126. 30Volts threshold power supply was ordered, as well.

Tests with first package at 126

- Beni found a way to dramatically reduce the noise and use thresholds <5Volts, compared to 13Volts before. Instead of the green wires, he used aluminum foil to connect directly the PCBs with the cathode grounds. Fernando and Beni also found some external noise.

- Fernando proposed different schemes for the grounding: using the holes in the periphery of the planes, or using clamps with copper braids. Beni and Fernando will investigate this on the first package and will come with a solution to be implemented on the third package; will need it in about a month from now. We discussed also how to reduce the material, but the first priority will be to have the chamber working at low threshold.